OpenStudy (anonymous):

15-28i=3l+(4m)i solve

OpenStudy (anonymous):

What are you solving for? Is i a variable or complex number?

OpenStudy (anonymous):

yes it is for complex number

OpenStudy (anonymous):

and that is a 3l like the letter L or is it a 31?

OpenStudy (anonymous):

it is the letter L, im trying to fine L & M

OpenStudy (anonymous):

For L 3L=15-28i-(4M)i 3L=15-4i(7+M) \[L=5-\frac{4i}{3}(7+M)\] For M (4M)i=15-28i-3L (4M)i=3(5-L)-28i (4M)=3(5-L)/i-28 (4M)=3(L-5)i-28 \[M=\frac{3(L-5)i}{4}-7\]

OpenStudy (anonymous):

Is this what you mean? :)

OpenStudy (anonymous):

not really, but thanks though :)